# Instruction Fusion: Advancing Prompt Evolution through Hybridization

### v2

@@ 1 Introduction
Addressing these challenges, we introduce a novel method named Instruction Fusion (IF). This method leverages hybridization concepts to significantly improve prompt evolution, as demonstrated in Figure 1. It involves merging two distinct instructions into a single prompt using GPT-4 Turbo, enhancing prompt complexity and diversity. This approach also facilitates a more gradual increase in difficulty for LLMs, optimizing training and performance.
To assess the efficacy of our Instruction Fusion method, we conducted experiments with Code-Llama Roziere et al. (2023), using HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), ${}^{+}$, ${}^{+}$ Liu et al. (2023a), and MultiPL-E Cassano et al. (2022) as benchmark datasets. Our results demonstrate a significant performance improvement of LLMs when training with the additional data generated by IF. The contributions of this work can be summarized as follows:
@@
  • Through extensive fine-tuning experiments on Code-Llama models on five commonly used code generation benchmarks, we demonstrate the superior performance of our Instruction Fusion method. We fully open source the model weights, training data, and source code at Anonymous to facilitate future research.
@@ 2 Approach
Figure 1: An illustration of Instruction Fusion: Pairs of evolved instructions are fused using prompts fed to GPT4-Turbo. Additionally, the responses to each fused instruction are also generated and collected from GPT4-Turbo.
In this section, we first analyze the the limitation of current evolution method on Code LLMs, then we we elaborate on the details of the proposed Instruction Fusing method and analyze its advantages comparing to the existing instruction evolution method. The main process of the fusion is illustrated in Figure 1.
@@ 2.1 Instruction Evolution
High-quality instructions are vital for effectively fine-tuning Code LLMs. Crafting these, particularly for coding tasks, is resource-intensive and often results in easy to moderate difficulty levels, creating a gap in challenging content Xu et al. (2023). To address this and reduce costs, Luo et al. (2023) applied Evol-Instruct Xu et al. (2023) to Code LLMs, enhancing code generation. This process involves merging each seed instruction with a unique evolution prompt and generating evolved instructions using GPT-3.5 Turbo. This method expanded and diversified the instruction dataset, but certain limitations hinder its further development.
@@
A major challenge identified in the process of instruction evolution is the escalation of constraints, as illustrated in Figure 2. The process begins with a straightforward instruction, such as developing a Python function to convert strings to lowercase. However, the complexity increases significantly with each evolution round. After four rounds, the instruction length expanded from 14 to 325 tokens. This increase is primarily attributed to the incorporation of up to eight distinct constraints. As a result, the evolved instructions become impractically intricate and overly complex.
Although Evol-Instruct aids in escalating difficulty levels by braining new constrains, which benefits LLM learning, those new constrains may need prior knowledge to learn. Consider math learning as an analogy: understanding addition and multiplication is fundamental before tackling problems like "$1+1\times 2$." However, Evol-Instruct might prematurely introduce such problems immediately following basic concepts like "$1+1=2$", potentially leading to gaps in learning progression.
@@ 2.2 Instruction Fusion
Here, $1\leq m\leq M$. For each fused instruction, if $I^{m}_{(j,k)}==\text{INVALID\ PROMPT}$, we will discard the $(j,k)$ combination and randomly sample a new one. We keep repeating the fusion process until the number of new instructions reach the target amount $M$. The result instructions set is then $\textbf{H}^{\text{IF}}=\{(I^{m}_{(j,k)},R^{m}_{(j,k)})\}_{1<m<M}$, where $R^{m}_{(j,k)}$ is the corresponding response of $I^{m}_{(j,k)}$ generated by GPT-4 Turbo.
Data Collection. Table 1 gives the statistic of the collected datasets. Due to the unavailability of WizardCoder’s Luo et al. (2023) original evolved instructions, we utilized the third-party dataset evol-codealpaca-v1111https://github.com/theblackcat102/evol-dataset, which mirrors the methods used by Luo et al. (2023). This dataset, comprising 111k refined instructions, is evolved from the CodeAlpaca [^2] https://github.com/sahil280114/codealpaca dataset, which is also serves as the seeds in WizardCoder. We divided evol-codealpaca-v1 based on programming language into Python-related instructions $\textbf{C}_{\text{PI}}$ and Non-Python-related instructions $\textbf{C}_{\text{NPI}}$, numbering 50,131 and 61,140, respectively. To perform detailed comparison in later experiments, we further divide $\textbf{C}_{\text{PI}}$ into $\textbf{C}_{30}$ and $\textbf{C}_{30r}$, which represents the 30K random samples from $\textbf{C}_{\text{PI}}$ and the rest samples, respectively. Based on the Instruction Fusion scheme introduced above, we set $\textbf{C}_{\text{PI}}$ as seeds and perform IF with $M=50,000$ to obtain the result dataset $\textbf{H}_{\text{PI}}$, where 20K fused instructions $\textbf{H}_{20}$ only use seeds from $\textbf{C}_{30}$. Then we perform another IF with $M=50,000$ on $\textbf{C}_{\text{PI}}+\textbf{C}_{\text{NPI}}$ and obtain $\textbf{H}_{\text{PI}+\text{NPI}}$. For cost efficiency, our method predominantly utilizes samples in $\textbf{H}_{\text{PI}}$. Nevertheless, towards the experiment’s end, we exam code generation performance using both two datasets, showcasing our method’s capability across multiple programming languages.
Table 1: Average tokens comparison. The token is obtained using the official tokenizer of CodeLlama.
Dataset # Instructions Inst. Avg. Tokens Resp. Avg. Tokens evol-codealpaca-v1 Python-only 50k 185.4 441.7 evol-codealpaca-v1 111k 209.3 438.9 $\mathbf{H}_{\text{PI}}$ 50k 222.4 712.0 $\mathbf{H}_{\text{PI+NPI}}$ 100k 260.4 754.8
@@ 2.3 Complement the Deficiencies
The effectiveness of this gradient is further substantiated by the theory presented in Kung et al. (2023)’s work. They categorize samples with high instruction uncertainty and a high prediction probability as “ambiguous”. Instruction uncertainty denotes the degree to which minor modifications in the instruction impact response generation. Higher uncertainty indicates greater sensitivity of LLMs to the specific instructions in a sample. Their research underscores the significant role of “ambiguous” training data in fine-tuning LLMs. Conversely, samples characterized by low instruction uncertainty and prediction probability are deemed challenging, potentially offering limited benefits for fine-tuning due to their complexity.
@@
Figure 5: tSNE plots of the sentiment embeddings of the instructions. The upper plot is for $\textbf{C}_{30}$ and $\textbf{C}_{30r}$. The lower plot represents $\textbf{C}_{30}$ and $\textbf{H}_{20}$.
Instruction Fusion (IF) counters the diversity limitations of Evol-Instruct, which is constrained by initial objectives, by creating instructions that blend the targets of parent instructions. This leads to unique tasks not found in the Evol-Instruct dataset. In Figure 5, the gray points represent a t-SNE 2D sentiment embedding plot for $\mathbf{C}_{30}$ instructions, showing small cliques indicative of limited diversity. The orange points represent samples from $\mathbf{C}_{30r}$. We can observe that the additional evolved data doesn’t fill the gap between those small cliques, which can demonstrate how the instruction diversity provided by Evol-Instruct is limited.
Conversely, the red points, representing IF’s fused instructions, populate previously empty areas, demonstrating increased diversity. The red points’ distribution fills gaps left by the blue points, illustrating a more uniform sentiment embedding.
@@
$\mu=\frac{1}{N}\Sigma^{N}_{i=1}(||(e_{k},e^{\text{NN}}_{k})||).$ (2)
Here, $U$ denotes distribution uniformity and $\mu$ is the average Euclidean Distance between the sentiment embedding $e_{k}$ of $I_{k}$ and the context embedding $e^{\text{NN}}_{k}$ of its nearest neighbor $I^{\text{NN}}_{k}$. Lower $U$ values suggest improved uniformity.
The results reveal notable findings. The first plot of $\textbf{C}_{\text{PI}}$, composed of $\textbf{C}_{30}+\textbf{C}_{30r}$, exhibits higher dispersion with a variance of 0.0332, indicating reduced uniformity compared to $\textbf{C}_{30}$ alone, which is 0.0316. However, an interesting contrast arises when employing Instruction Fusion on $\textbf{C}_{30}$: by integrating $\textbf{H}_{20}$ with $\textbf{C}_{30}$, the dispersion of the second plot decreases by 28.5% from 0.0332 to 0.0226. This significant reduction suggests that Instruction Fusion effectively narrows the gaps in sentiment embeddings between existing instructions. This improvement is achieved by combining the objectives of two distinct instructions, demonstrating the method’s efficacy in enhancing instruction diversity.
In conclusion, the Instruction Fusion (IF) method we propose serves as a valuable complement to Evol-Instruct, effectively addressing its limitations. By amalgamating diverse objectives from various seed instructions, IF leads to the creation of a more challenging and diverse set of instructions. This method not only expands the range of tasks addressed but also ensures a more harmonious and gradual escalation in the complexity of instructions.
3 Experiments
Table 2: Results of experiments on Python benchmarks. The abbreviations ’CL’ and ’CLP’ denote the base models CodeLlama and CodeLlama-Python, respectively. ’EC’ and ’ECP’ represent the models fine-tuned with evol-codealpaca-v1 and its Python-only variant. The values in bracket indicate the performance improvement achieved by fine-tuning with instructions generated by IF, as compared to the use of evol-codealpaca-v1 alone.
@@ 3 Experiments
IF-CLP-ECP 13B weight&data 73.8 69.5 71.7 61.7
IF-CL-ECP 13B weight&data 75.6 68.9 67.9 58.6
IF-CL-EC 13B weight&data 74.4 68.3 69.7 59.4
IF-CLP-ECP 34B weight&data 75.6 69.5 73.7 62.7
IF-CL-ECP 34B weight&data 76.8 70.7 73.2 62.9
IF-CL-EC 34B weight&data 78.7 71.3 71.4 60.7
In this section, we report the experiment details of LLMs fine-tuned with instruction generated by IF. We focused on the pass@1 performance under greedy generation settings across five benchmarks: four Python benchmarks (HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), ${}^{+}$, ${}^{+}$ Liu et al. (2023a)) and the multi-language benchmark MultiPL-E Cassano et al. (2022). The official EvalPlus code [^3] https://github.com/evalplus/evalplus was used for evaluating the Python benchmarks, while bigcode-evaluation-harness Ben Allal et al. (2022) was utilized to evaluate the performance on MultiPL-E.
@@ 3.1 Experiment Details
To test Instruction Fusion, we fine-tune both the basic and Python versions of CodeLlama 13 billion (13B) [^6] https://huggingface.co/codellama/CodeLlama-13b-hf and 34 billion (34B) [^7] https://huggingface.co/codellama/CodeLlama-34b-hf as base models. For fine-tuning, models were trained with a batch size of 256 over 2 epochs, a learning rate of 2e-5, a cosine learning rate scheduler with 10% warm-up steps, and under bf16 precision on $4\times 8$ NVIDIA A100 40G GPUs. All other hyper-parameters remained consistent with WizardCoder and Magicoder unless specified otherwise.
3.2 Evaluation
Table 2 presents the performance of various open-source models on Python benchmarks. When compared to the leading closed-source models, GPT-3.5 Turbo, our models with 13B and 34B parameters demonstrate superior performance on the HumanEval and ${}^{+}$ benchmarks. However, there is a noticeable performance gap on the two MPBB benchmarks. This discrepancy may stem from the MBPP test cases, which often require specific prior knowledge for resolution. For instance, in the real test case task_id:20, the prompt "Write a function to check if a given number is woodball or not." requires the model’s understanding of what constitutes a ’woodball’ number. This type of prior knowledge is either not acquired during the pre-training stage or forgotten during the supervised fine-tuning process, leading to a consistent failure across all IF models for this case.
However, compared with all open-source models, the IF models excel in Python code generation performance, surpassing all baselines. The fused instructions, which offer richer sentiment, increased difficulty, and smoother transitions, allow both the 13B and 34B versions of IF models to achieve new benchmarks, outperforming models of comparable sizes. Notably, the 13B version of IF-CLP-ECP outperforms all 34B baselines in the HumanEval and ${}^{+}$ benchmarks and shows comparable results on MBPP and ${}^{+}$.
@@ 3.3 Ablation Study
Method HumanEval HumanEval+ MBPP MBPP+ IF-CLP-ECP 73.8 69.5 71.7 61.7   -fused inst. 67.7 64.0 66.4 56.4 IF-CL-ECP 75.6 68.9 67.9 58.6   -fused inst. 66.5 61.6 67.2 57.4 IF-CL-EC 74.4 68.3 69.7 59.4   -fused inst. 65.2 61.0 68.4 56.1 IF-CLP-ECP 75.6 69.5 73.7 62.7   -fused inst. 72.0 65.2 72.2 61.9 IF-CL-ECP 76.8 70.7 73.2 62.9   -fused inst. 67.7 62.8 70.9 61.2 IF-CL-EC 78.7 71.3 71.4 60.7   -fused inst. 67.7 62.2 69.9 60.4

### v4

@@ 1 Introduction
Addressing these challenges, we introduce a novel method named Instruction Fusion (IF). This method leverages hybridization concepts to significantly improve prompt evolution, as demonstrated in Figure 1. It involves merging two distinct instructions into a single prompt using GPT-4 Turbo, enhancing prompt complexity and diversity. This approach also facilitates a more gradual increase in difficulty for LLMs, optimizing learning and performance.
To assess the efficacy of our Instruction Fusion method, we conducted experiments with Code-Llama Roziere et al. (2023), using HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), HumanEval+, MBPP+ Liu et al. (2023a), and MultiPL-E Cassano et al. (2022) as benchmark datasets. Our results demonstrate a significant performance improvement of LLMs when training with the additional data generated by IF. The contributions of this work can be summarized as follows:
@@
  • Through extensive fine-tuning experiments on Code-Llama models on five commonly used code generation benchmarks, we demonstrate the superior performance of our Instruction Fusion method. We fully open source the model weights, training data, and source code to facilitate future research.
Figure 1: An illustration of Instruction Fusion. The responses to each fused instruction are also generated and collected from GPT-4 Turbo.
@@ 2 Approach
In this section, we first analyze the the limitation of current evolution method on Code LLMs, then we elaborate on the details of the proposed Instruction Fusion method and analyze its advantages comparing to the existing instruction evolution method. The main process of the fusion is illustrated in Figure 1.
@@ 2.1 Instruction Evolution
High-quality instructions are vital for effectively fine-tuning Code LLMs. Crafting these, particularly for coding tasks, is resource-intensive and often results in easy difficulty levels, creating a gap in challenging content Xu et al. (2023). To address this and reduce costs, Luo et al. (2023) applied Evol-Instruct Xu et al. (2023) to Code LLMs, enhancing code generation. This process involves merging each seed instruction with a unique evolution prompt and generating evolved instructions using GPT-3.5 Turbo. This method expanded and diversified the instruction dataset, but certain limitations hinder its further development.
@@
A major challenge identified in the process of instruction evolution is the over-escalating constraints, as illustrated in Figure 2. The process begins with a straightforward instruction, such as developing a Python function to convert strings to lowercase. However, the complexity increases significantly with each evolution round. After four rounds, the instruction length expanded from 14 to 325 tokens. This increase is primarily attributed to the incorporation of up to eight distinct constraints. As a result, the evolved instructions become impractically intricate and overly complex.
Although Evol-Instruct with low rounds can properly increase the difficulty levels by braining new constrains, which benefits LLM learning, those new constrains may need prior knowledge to learn. Consider math learning as an analogy: understanding addition and multiplication is fundamental before tackling problems like "$1+1\times 2$." However, Evol-Instruct might prematurely introduce such problems immediately following basic concepts like "$1+1=2$", potentially leading to gaps in learning progression.
@@ 2.2 Instruction Fusion
Here, $1\leq m\leq M$. For each fused instruction, if $I^{m}_{(j,k)}==\text{INVALID\ PROMPT}$, we will discard the $(j,k)$ combination and randomly sample a new one. We keep repeating the fusion process until the number of new instructions reach the target amount $M$. The result instructions set is then $\textbf{H}=\{(I^{m}_{(j,k)},R^{m}_{(j,k)})\}_{1<m<M}$, where $R^{m}_{(j,k)}$ is the corresponding response of $I^{m}_{(j,k)}$ generated by GPT-4 Turbo (gpt-4-1106-preview [^1] https://platform.openai.com/docs/models).
Data Collection. Table 1 gives the statistic of the collected datasets. Due to the unavailability of WizardCoder’s Luo et al. (2023) original evolved instructions, we utilized the third-party dataset evol-codealpaca-v1222https://github.com/theblackcat102/evol-dataset, which mirrors the methods used by Luo et al. (2023). This dataset, comprising 111k refined instructions, is evolved from the CodeAlpaca [^3] https://github.com/sahil280114/codealpaca dataset, which is also serves as the seeds in WizardCoder. We divided evol-codealpaca-v1 based on programming language into Python-related instructions $\textbf{C}_{\text{PI}}$ and Non-Python-related instructions $\textbf{C}_{\text{NPI}}$, numbering 50,131 and 61,140, respectively. To perform detailed comparison in later experiments, we further divide $\textbf{C}_{\text{PI}}$ into $\textbf{C}_{30}$ and $\textbf{C}_{30r}$, which represents the 30K random samples from $\textbf{C}_{\text{PI}}$ and the rest samples, respectively. Based on the Instruction Fusion scheme introduced above, we set $\textbf{C}_{\text{PI}}$ as seeds and perform IF with $M=50,000$ to obtain the result dataset $\textbf{H}_{\text{PI}}$, where 20K fused instructions $\textbf{H}_{20}$ only use seeds from $\textbf{C}_{30}$. Then we perform another IF with $M=50,000$ on $\textbf{C}_{\text{PI}}+\textbf{C}_{\text{NPI}}$ and obtain $\textbf{H}_{\text{PI}+\text{NPI}}$. For cost efficiency, our method predominantly utilizes samples in $\textbf{H}_{\text{PI}}$. Nevertheless, towards the experiment’s end, we exam code generation performance using both two datasets, showcasing our method’s capability across multiple programming languages.
Table 1: Average tokens comparison. The token is obtained using the official tokenizer of CodeLlama. ‘EC’ denotes evol-codealpaca-v1, while ‘ECP’ refers to its Python-only variant.
Dataset # Instructions Inst. Avg. Tokens Resp. Avg. Tokens ECP 50k 185.4 441.7 EC 111k 209.3 438.9 $\mathbf{H}_{\text{PI}}$ 50k 222.4 712.0 $\mathbf{H}_{\text{PI+NPI}}$ 100k 260.4 754.8
@@ 2.3 Complement the Deficiencies
The effectiveness of this gradient is further substantiated by the theory presented in Kung et al. (2023)’s work. They categorize samples with high instruction uncertainty and a high prediction probability as “ambiguous”. Instruction uncertainty denotes the degree to which minor modifications in the instruction impact response generation. Higher uncertainty indicates greater sensitivity of LLMs to the specific instructions in a sample. Their research underscores the significant role of “ambiguous” training data in fine-tuning LLMs. Conversely, samples characterized by low instruction uncertainty and prediction probability are deemed over-challenging, potentially offering limited benefits for fine-tuning due to their complexity.
@@
Figure 5: tSNE plots of the semantic embeddings of the instructions. The upper plot is for $\textbf{C}_{30}$ and $\textbf{C}_{30r}$. The lower plot represents $\textbf{C}_{30}$ and $\textbf{H}_{20}$.
Instruction Fusion (IF) counters the diversity limitations of Evol-Instruct, which is constrained by initial objectives, by creating instructions that blend the targets of parent instructions. This leads to unique tasks not found in the Evol-Instruct dataset. In Figure 5, the gray points represent a t-SNE 2D semantic embedding plot for $\mathbf{C}_{30}$ instructions, The upper plot represent $\mathbf{C}_{\text{PI}}$, showing small cliques indicative of limited diversity, and the orange points represent samples from $\mathbf{C}_{30r}$.
In the lower plot, the red points, representing IF’s fused instructions, populate previously empty areas, demonstrating increased diversity, and illustrating a more uniform semantic embedding.
@@
$d_{i}=||(e_{i},e^{\text{NN}}_{i})||$ (2)
and
$\mu=\frac{1}{N}\Sigma^{N}_{i=1}d_{i}.$ (3)
Here, $U$ denotes distribution uniformity, $d_{i}$ is the Euclidean Distance between the semantic embedding $e_{i}$ of $I_{i}$ and the context embedding $e^{\text{NN}}_{i}$ of its nearest neighbor $I^{\text{NN}}_{i}$, while $\mu$ is the average Euclidean Distance. Lower $U$ values suggest improved uniformity. In ideal case, if all points are distributed uniformly, their distance to their nearest neighbor should be equal, and the variance should be zero. In other words, lower the variance indicates more uniform semantic embedding among instructions.
The results reveal notable findings. The first plot of $\textbf{C}_{\text{PI}}$, composed of $\textbf{C}_{30}+\textbf{C}_{30r}$, exhibits higher dispersion with a variance of 0.0332, indicating reduced uniformity compared to $\textbf{C}_{30}$ alone, which is 0.0316. However, an interesting contrast arises when employing Instruction Fusion on $\textbf{C}_{30}$: by integrating $\textbf{H}_{20}$ with $\textbf{C}_{30}$, the dispersion of the second plot decreases by 28.5% from 0.0316 to 0.0226. This significant reduction suggests that Instruction Fusion effectively narrows the gaps in semantic embeddings between existing instructions. This improvement is achieved by combining the objectives of two distinct instructions, demonstrating the method’s efficacy in enhancing instruction diversity.
In conclusion, the Instruction Fusion (IF) method we propose serves as a valuable complement to Evol-Instruct, effectively addressing its limitations. By amalgamating diverse objectives from various seed instructions, IF leads to the creation of a more diverse, challenging instructions with gradual escalation in the complexity.
Table 2: Results of experiments on Python benchmarks. The abbreviations ‘CL’ and ‘CLP’ denote the base models CodeLlama and CodeLlama-Python, respectively. For IF method, we fine-tune CodeLlama with full evol-codealpaca-v1 dataset and CodeLlama-Python with only Python instruction in evol-codealpaca-v1.
@@
IF-CLP 13B weight&data 73.8 69.5 71.7 61.7
IF-CL 13B weight&data 74.4 68.3 69.7 59.4
IF-CLP 34B weight&data 75.6 69.5 73.7 62.7
IF-CL 34B weight&data 78.7 71.3 71.4 60.7
3 Experiments
In this section, we report the experiment details of LLMs fine-tuned with instruction generated by IF. We focused on the pass@1 performance under greedy generation settings across five benchmarks: four Python benchmarks (HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), HumanEval+, MBPP+ Liu et al. (2023a)) and the multi-language benchmark MultiPL-E Cassano et al. (2022). The official EvalPlus code [^4] https://github.com/evalplus/evalplus was used for evaluating the Python benchmarks, while bigcode-evaluation-harness Ben Allal et al. (2022) was utilized to evaluate the performance on MultiPL-E.
@@ 3.1 Experiment Details
To test Instruction Fusion, we fine-tune both the basic and Python versions of CodeLlama 13 billion (13B) and 34 billion (34B) [^7] https://huggingface.co/codellama/ as base models. For fine-tuning, models were trained with a batch size of 256 over 2 epochs, a learning rate of 2e-5, a cosine learning rate scheduler with 10% warm-up steps, and under bf16 precision on $4\times 8$ NVIDIA A100 40G GPUs. All other hyper-parameters remained consistent with WizardCoder and Magicoder unless specified otherwise.
@@
3.2 Evaluation
Table 2 presents the performance of various open-source models on Python benchmarks. When compared to the leading closed-source models, GPT-3.5 Turbo, our models with 13B and 34B parameters demonstrate superior performance on the HumanEval and HumanEval+ benchmarks. However, there is a noticeable performance gap on the two MPBB benchmarks. This discrepancy may stem from the MBPP test cases, which often require specific prior knowledge for resolution. For instance, in the real test case task_id:20, the prompt "Write a function to check if a given number is woodball or not." requires the model’s understanding of what constitutes a ’woodball’ number. This type of prior knowledge is either not acquired during the pre-training stage or forgotten during the supervised fine-tuning process, leading to a consistent failure across all IF models for this case.
However, compared with all open-source models, the IF models excel in Python code generation performance, surpassing all baselines. The fused instructions, which offer richer semantic, increased difficulty, and smoother transitions, allow both the 13B and 34B versions of IF models to achieve new state-of-the-art performance, outperforming models of comparable sizes. Notably, the 13B version of IF-CLP outperforms all 34B baselines in the HumanEval and HumanEval+ benchmarks and shows comparable results on MBPP and MBPP+.
@@ 3.3 Ablation Study
Method HumanEval HumanEval+ MBPP MBPP+ IF-CLP 73.8 69.5 71.7 61.7     -fused inst. 67.7 64.0 66.4 56.4 IF-CL 74.4 68.3 69.7 59.4     -fused inst. 65.2 61.0 68.4 56.1 IF-CLP 75.6 69.5 73.7 62.7     -fused inst. 72.0 65.2 72.2 61.9 IF-CL 78.7 71.3 71.4 60.7     -fused inst. 67.7 62.2 69.9 60.4